OpenStudy (anonymous):

Find the slope-intercept form for the line that satisfies the following conditions: x-intercept 3, y-intercept 3/5

OpenStudy (anikay):

y=mx+b b= 3/5

OpenStudy (anikay):

so m=1/5

OpenStudy (anikay):

y=x/5+3/5

OpenStudy (anikay):

or y= (x+3)/5

OpenStudy (anonymous):

Ohh Ok.. Thank you.